Geometry. The length of a rectangle is 2 in. more than twice its width. If the
perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by checkley75(3666)

LENGTH=2W+2
THE PERIMETER IS 2L+2W THUS:
2(2W+2)+2W=34
4W+4+2W=34
6W=34-4
6W=30
W=30/6
W=5 ANSWER.
L=2*5+2
L=10+2
L=12 ANSWER.
PROOF
2*5+2*12=34
10+24=34
34=34